# Semi-Blind Strategies for MMSE Channel Estimation Utilizing Generative Priors

### v1

@@ II-B Measurement Campaign
Since synthetic data capture real-world CSI characteristics only up to some extent, we utilize real-world data from a measurement campaign conducted at the Nokia campus in Stuttgart, Germany, in October/November 2017, cf. [36]. The BS antenna with a uniform rectangular array (URA) comprises $N_{v}=4$ vertical ($\lambda$ spacing) and $N_{h}=16$ horizontal ($\lambda/2$ spacing) single polarized patch antennas. The operating carrier frequency is $2.18$ GHz and the antenna was mounted on a rooftop approximately $20$ meters above the ground. For further details, we refer the reader to [36].
@@ III Semi-Blind Channel Estimation using Perfect Statistical Knowledge
In this section, we introduce two variants of the LMMSE estimator incorporating information provided by the payload data symbols. To do so, we first restrict ourselves to the case where perfect statistical knowledge is available at the receiver.
In channel estimation, commonly, only the pilot observation ${\bm{y}}_{p}$ is considered for channel estimation. The MSE optimal estimator given the pilot observation ${\bm{y}}_{p}$ is the CME
@@
where $\rho_{i}$ are the eigenvalues of ${\bm{C}}$.
For our considerations concerning semi-blind channel estimation, we assume knowledge about the subspace defined by $\mathrm{range}({\bm{H}})=\mathrm{range}({\bm{V}})$, where we denote with ${\bm{V}}$ the left singular vectors of ${\bm{H}}$, which span the same subspace as the columns of ${\bm{H}}$. Generally, we can formulate the ML estimate of ${\bm{H}}$ in view of (2) as [16]
@@
where the first term belongs to the pilot observations and the second part refers to the observation obtained from the payload symbols. In [16] the EM algorithm is introduced to solve this channel estimation problem in terms of maximum likelihood, whereas, in [11] a semi-blind method is derived based on utilizing the subspace $\mathrm{range}({\bm{H}})=\mathrm{range}({\bm{V}})$. Now, given the subspace $\mathrm{range}({\bm{V}})$, we can reformulate (13) as [11]
$\min_{{\bm{S}},{\bm{X}}}\sum_{n=1}^{J}\left\|{\bm{y}}(n)-{\bm{V}}{\bm{s}}_{n}\right\|^{2}+\sum_{n=J+1}^{N}\left\|{\bm{y}}(n)-{\bm{V}}{\bm{S}}{\bm{x}}(n)\right\|^{2},$ (14)
with ${\bm{H}}={\bm{V}}{\bm{S}}$ and ${\bm{S}}=[{\bm{s}}_{1},...,{\bm{s}}_{J}]\in\mathbb{C}^{J\times J}$. The optimal solution for ${\bm{X}}$ is given as
${\bm{x}}^{*}(n)$ $=({\bm{S}}^{\mathrm{H}}{\bm{V}}^{\mathrm{H}}{\bm{V}}{\bm{S}})^{-1}{\bm{S}}^{\mathrm{H}}{\bm{V}}^{\mathrm{H}}{\bm{y}}(n)$ (15)
$=({\bm{S}}^{\mathrm{H}}{\bm{S}})^{-1}{\bm{S}}^{\mathrm{H}}{\bm{V}}^{\mathrm{H}}{\bm{y}}(n).$ (16)
Reinserting this solution into (13), the right part of the objective simplifies to
$\min_{{\bm{S}}}\sum_{n=1}^{J}\left\|{\bm{y}}(n)-{\bm{V}}{\bm{s}}_{n}\right\|^{2}+\sum_{n=J+1}^{N}\left\|{\bm{y}}(n)-{\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{y}}(n)\right\|^{2},$ (17)
where only the left part depends on ${\bm{S}}$. Thus, the ML problem for the user of interest results in [11]
@@
with the closed form solution $\hat{{\bm{h}}}_{\text{ML}}={\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{y}}_{p}$. The MSE of this estimator is
@@ III-A Subspace Channel Estimator
Using the information in $\mathrm{range}({{\bm{V}}})$, we can solve the estimation within the subspace as previously proposed in [26]. For this, the pilot system model in (4) is projected into the $J$-dimensional subspace as
@@
One should note that by design ${\bm{V}}$ actually depends on ${\bm{h}}$ and, hence, (21) is a suboptimal but feasible estimate for ${\bm{h}}^{\prime}$. After solving the estimation in the subspace for ${\bm{h}}^{\prime}$, the solution can be transformed back using [26]
@@ III-B Projected Channel Estimator
As an alternative approach, we propose the use of the orthogonal subspace projection ${\bm{P}}_{\bm{H}}={\bm{V}}{\bm{V}}^{\mathrm{H}}$ as a preprocessing filter. Since the projector ${\bm{P}}_{\bm{H}}$ does not affect ${\bm{h}}$, the resulting projected observation is given by
@@
To formulate the CME $\hat{{\bm{h}}}=\mathbb{E}\left[{\bm{h}}\mid\tilde{{\bm{y}}}\right]$, we need to calculate the statistic of the noise $\tilde{{\bm{n}}}$ with
${\bm{C}}_{\tilde{{\bm{n}}}}=\;$ $\mathbb{E}\left[\tilde{{\bm{n}}}\tilde{{\bm{n}}}^{\mathrm{H}}\right]=\mathbb{E}\left[\sigma^{2}{\bm{P}}_{\bm{H}}\right].$ (24)
To get an intuitive understanding of (24), let us consider a scenario involving spatially uncorrelated channels, meaning that path gains and channel directions are uncorrelated. This is the case when users are uniformly distributed over the directions, e.g., the spatial channel model in Section II-A, resulting in a channel covariance matrix of the scenario that is a scaled identity [37, Def. 2.3]. In such a case, the matrices with the eigenvectors of the sample covariance matrix of such channels are distributed with Haar measure [38, Chap. 1], i.e., uniformly distributed on the manifold of unitary matrices. Assuming spatially uncorrelated channels (24) results in
@@
III-C Performance Analysis
@@
Comparing the performance to the plain LMMSE we see that
@@
holds for every $i=1,\dots,M$, resulting in $\mathrm{MSE}^{\mathrm{proj}}\leq\mathrm{MSE}^{\mathrm{plain}}$, cf. (12). The inequality in (29) only holds with equality if $J=M$. Additionally, we can compare the MSE of the projected LMMSE to (19) by reformulating (28) as
@@
To compare the projected variant to the subspace LMMSE, let us consider uncorrelated Rayleigh fading such that the channel covariance matrix is given as ${\bm{C}}=\mathbf{I}_{M}$. The MSE of the projected variant results in
@@
For the case of ${\bm{C}}=\mathbf{I}_{M}$, the subspace LMMSE estimator boils down to
@@
Fig. 1 shows the performance of the individual channel estimators based on perfect genie knowledge for the case of uncorrelated Rayleigh fading. As one can see, the projected channel estimator outperforms all other estimators across the whole signal-to-noise ratio (SNR) range. Further, we realize that the subspace LMMSE converges to the ML method from above.
@@
Figure 1: MSE over the SNR for given channel estimations based on perfect subspace and perfect statistical knowledge in a $J=8$ user and $M=64$ antennas scenario with uncorrelated Rayleigh fading.
@@ IV Proposed Semi-Blind Channel Estimation - Utilizing Generative Prior
In practice, (10) can not be utilized directly as the channels have to be Gaussian distributed, and this distribution needs to be known. In general, the CME given the pilot observation ${\bm{y}}_{p}$ is formulated as
@@
As can be seen in (34), the CME generally can not be computed analytically. First, the CME needs access to $p({\bm{h}})$, which is generally unavailable in practice. Additionally, no closed-form solution exists to the integral in (34).
@@
denotes the LMMSE estimate given ${\bm{c}}$. However, finding a suitable condition $\bm{\delta}$ can be challenging, in particular as the true distribution of ${\bm{h}}$ is unknown. To this end, CGLMs were proposed in [30, 31, 32], which approximate the CME based on a GMM, mixtures of factor analyzers (MFA), and VAE, respectively. All three methods learn a model that provides the conditional Gaussian distribution ${\bm{h}}\mid{\bm{c}}$ based on a discrete (GMM, MFA) or continuous (VAE) latent variable ${\bm{c}}$. In this work, we focus on the GMM and VAE, which we adapt to semi-blind channel estimation in the following.

### v2

@@ II-B Measurement Campaign
Since synthetic data captures real-world CSI characteristics only up to some extent, we utilize real-world data from a measurement campaign conducted at the Nokia campus in Stuttgart, Germany, in October/November 2017, cf. [36]. The BS antenna with a uniform rectangular array (URA) comprises $N_{v}=4$ vertical ($\lambda$ spacing) and $N_{h}=16$ horizontal ($\lambda/2$ spacing) single polarized patch antennas. The operating carrier frequency is $2.18$ GHz, and the antenna was mounted on a rooftop approximately $20$ meters above the ground. For further details, we refer the reader to [36].
@@ III Semi-Blind Channel Estimation using Perfect Statistical Knowledge
In this section, we introduce two variants of the LMMSE estimator incorporating information provided by the payload data symbols. To do so, we restrict ourselves for now to the case where perfect statistical knowledge is available at the receiver.
In channel estimation, commonly, only the pilot observation ${\bm{y}}_{p}$ is considered for channel estimation. The MSE-optimal estimator given the pilot observation ${\bm{y}}_{p}$ is the CME
@@
where $\rho_{i}$ are the eigenvalues of $\bm{C}$}. In contrast to MMSE-based channel estimation, ML-based channel estimation does not utilize knowledge about the channels’ prior distribution, making it a suboptimal estimation problem.
For our considerations concerning semi-blind channel estimation, we assume knowledge about the subspace defined by $\mathrm{range}({\bm{H}})=\mathrm{range}({\bm{V}})$, where we denote with $\bm{V}$} the left singular vectors of $\bm{H}$}, which span the same subspace as the columns of $\bm{H}$}. Generally, we can formulate the (suboptimal) ML estimate of $\bm{H}$} in view of (2) as [15]
@@
where the first term belongs to the pilot observations, and the second part refers to the observation obtained from the payload symbols. Obtaining a closed-form solution for (13) is known to be hard [37]. In [15], the EM algorithm is introduced to solve this channel estimation problem in terms of maximum likelihood, whereas, in [11], a semi-blind method is derived based on utilizing the subspace $\mathrm{range}({\bm{H}})=\mathrm{range}({\bm{V}})$. Now, given the subspace $\mathrm{range}({\bm{V}})$, we can reformulate (13) as [11]
$\min_{{\bm{S}},{\bm{X}}}\sum_{n=1}^{J}\left\|{\bm{y}}(n)-{\bm{V}}{\bm{s}}_{n}\right\|^{2}+\hskip-1.0pt\sum_{n=J+1}^{N}\hskip-1.0pt\left\|{\bm{y}}(n)-{\bm{V}}{\bm{S}}{\bm{x}}(n)\right\|^{2},$ (14)
with ${\bm{H}}={\bm{V}}{\bm{S}}$ and ${\bm{S}}=[{\bm{s}}_{1},\dots,{\bm{s}}_{J}]\in\mathbb{C}^{J\times J}$. We can rewrite the second term in (14) as
$\sum_{n=J+1}^{N}\hskip-5.0pt\left\|{\bm{y}}(n)-{\bm{V}}{\bm{S}}{\bm{x}}(n)\right\|^{2}$ $=\hskip-5.0pt\sum_{n=J+1}^{N}\hskip-5.0pt\left\|\left(\mathbf{I}_{M}-{\bm{V}}{\bm{V}}^{\mathrm{H}}\right){\bm{y}}(n)\right\|^{2}$
$\quad+\left\|{\bm{V}}^{\mathrm{H}}{\bm{y}}(n)-{\bm{S}}{\bm{x}}(n)\right\|^{2},$ (15)
where the equality comes from the fact that the two terms are orthogonal to each other. The first term in (15) is constant for given $\bm{V}$}. In the case of a continuous constellation for ${\bm{x}}(n)$, the second term vanishes by solving for $\bm{X}$}. Further, in the case of discrete symmetric signal constellations, the second term in (15) can not be jointly solved for $\bm{S}$} and $\bm{X}$} due to phase ambiguity. Thus, the ML problem for the user of interest reduces to the pilot observations as [11]
@@
with the closed form solution $\hat{{\bm{h}}}_{\text{ML}}={\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{y}}_{p}$. This estimator’s MSE is
@@ III-A Subspace Channel Estimator
Using the information in $\mathrm{range}({{\bm{V}}})$, we can solve the estimation within the subspace as previously proposed in [26]. For this, the pilot system model in (4) is mapped into the $J$-dimensional subspace as
@@
By design, $\bm{V}$} depends on $\bm{h}$} and, hence, (19) is a suboptimal but feasible estimate for $\bm{h}$$^{\prime}$. After solving the estimation in the subspace for $\bm{h}$$^{\prime}$, the solution can be transformed back using [26]
@@ III-B Projected Channel Estimator
As an alternative approach, we propose using the orthogonal subspace projection ${\bm{P}}_{\bm{H}}={\bm{V}}{\bm{V}}^{\mathrm{H}}$ as a preprocessing filter. Since the projector ${\bm{P}}_{\bm{H}}$ does not affect $\bm{h}$}, the resulting projected pilot observation is given by
@@
To formulate the CME $\hat{{\bm{h}}}=\mathbb{E}\left[{\bm{h}}\mid\tilde{{\bm{y}}}\right$], we calculate the covariance matrix of $\tilde{{\bm{y}}}$ as
$\mathbb{E}\left[\tilde{{\bm{y}}}\tilde{{\bm{y}}}^{\mathrm{H}}\right]$ $={\bm{C}}+\mathbb{E}\left[\sigma^{2}{\bm{P}}_{\bm{H}}\right],$ (22)
where the mixing terms vanish due to the independence of $\bm{h}$} and $\bm{n}$}. To get an intuitive understanding of the noise covariance in (22), let us consider a scenario involving spatially uncorrelated channels, meaning that path gains and channel directions are uncorrelated. This is the case when users are uniformly distributed over the directions, e.g., the spatial channel model in Section II-A, resulting in a scenario’s channel covariance matrix that is a scaled identity [38, Def. 2.3]. In such a case, the matrices with the sample covariance matrix’s eigenvectors are distributed with Haar measure [39, Chap. 1], i.e., uniformly distributed on the manifold of unitary matrices. Assuming spatially uncorrelated channels we have
@@
III-C Performance Analysis for Perfect Subspace Knowledge
@@
Comparing the performance to the plain LMMSE from (12) we see that
@@
holds for every $i\in\{1,\dots,M\}$, resulting in $\mathrm{MSE}^{\mathrm{proj}}\leq\mathrm{MSE}^{\mathrm{plain}}$. The inequality in (27) only holds with equality if $J=M$. Additionally, we can compare the MSE of the projected LMMSE to the ML estimator from (17) by reformulating (26) as
@@
In the case of uncorrelated Rayleigh fading, the channel covariance matrix is given as ${\bm{C}}=\mathbf{I}_{M}$. The projected variant’s MSE results in
@@
Now let us compare the projected variant to the subspace LMMSE. For the case of ${\bm{C}}=\mathbf{I}_{M}$, the subspace LMMSE estimator boils down to
@@
Fig. 1a shows the individual channel estimators’ performances based on perfect subspace knowledge for the case of uncorrelated Rayleigh fading. As one can see, the projected channel estimator outperforms all other estimators across the whole SNR range. Further, we realize that the subspace LMMSE converges to the ML method from above.
Additionally, performance guarantees of the proposed projected LMMSE compared to the unbiased CRB as derived in [15] can be given. We formulate the following theorem for the asymptotic region, where $M$ and $N$ tend to infinity.
Theorem 1.
We consider the limits of $N\to\infty$ and $M/N\to\alpha$ with $\alpha>0$, while keeping the number of pilot symbols in the order of users $J$. Then, the projected LMMSE based on perfect knowledge of the subspace spanned by the left singular vectors of the channel matrix $\bm{H}$} is equal to or better than any unbiased estimator in terms of MSE for uncorrelated Rayleigh fading.
Proof.
We take the limit of the MSE of the projected LMMSE estimator as
$\lim_{\begin{subarray}{c}N\to\infty\\ \frac{M}{N}\to\alpha\end{subarray}}\mathrm{MSE}^{\mathrm{proj}}$ $=\lim_{\begin{subarray}{c}N\to\infty\\ \frac{M}{N}\to\alpha\end{subarray}}\frac{J}{M}\sigma^{2}\sum_{i=1}^{M}\frac{\rho_{i}}{\rho_{i}+\frac{J}{M}\sigma^{2}}$ (32)
$=J\sigma^{2}\leq(1+\alpha)J\sigma^{2}=\mathrm{CRB}_{\mathrm{iid,d}},$ (33)
where $\mathrm{CRB}_{\mathrm{iid,d}}$ denotes the deterministic CRB’s limit for uncorrelated Rayleigh fading. The last equality in (33) was proven in [40, Theorem 4]. Furthermore, it is stated in [40] that in this asymptotic limit, the deterministic CRB is always lower than the stochastic CRB. ∎
Theorem 1 provides some insights on the asymptotic superiority of our proposed projected LMMSE strategy based on perfect subspace knowledge compared to any unbiased channel estimator. Additionally, the MSE of the subspace variant as well as the plain LMMSE approach infinity for the considered limit, making them irrelevant for this asymptotic regime. Interestingly, the projected LMMSE also approaches $J\sigma^{2}$ for the case where only $M$ tends to infinity, making the projected LMMSE a valid candidate for large antenna systems as the performance no longer depends on $M$.
III-D Maximum Likelihood Subspace Estimation
After introducing the methods utilizing the additional subspace information provided by $\mathrm{range}({\bm{V}})$ to enhance the CSI estimation quality, let us consider the estimation of such a subspace. As the received data symbols are transmitted over the same channels, we can use these payload symbols to estimate the subspace containing all user channels.
To this end, let us reconsider the ML estimate of $\bm{H}$} in (13). Instead of directly optimizing on this ML formulation as done in [15], which generally does not result in the MMSE, we only take this log-likelihood formulation as an intermediate step to estimate the subspace $\mathrm{range}({\bm{V}})$. First, let us consider the payload ML objective formulation in (15). We can then reformulate the problem, neglecting the second term due to phase ambiguity, resulting in [11]
where ${\bm{P}}_{\bm{H}}={\bm{H}}({\bm{H}}^{\mathrm{H}}{\bm{H}})^{-1}{\bm{H}}^{\mathrm{H}}={\bm{V}}{\bm{V}}^{\mathrm{H}}$ and $\hat{{\bm{C}}}^{(d)}_{{\bm{y}}\mid{\bm{H}}}=\frac{1}{N-J}{\bm{Y}}_{d}{\bm{Y}}_{d}^{\mathrm{H}}$, with ${\bm{Y}}_{d}$ from (2). The maximization in (34) is solved by setting ${\bm{P}}_{\bm{H}}$ equal to $\hat{{\bm{V}}}_{d}\hat{{\bm{V}}}_{d}^{\mathrm{H}}$ with $\hat{{\bm{V}}}_{d}$ holding the $J$ dominant eigenvectors of the receive sample covariance matrix $\hat{{\bm{C}}}^{(d)}_{{\bm{y}}\mid{\bm{H}}}$. Additionally, it is trivial to see that the first term in (13) is minimized by ${\bm{h}}_{n}={\bm{y}}(n)$. The subspace spanned by the solution ${\bm{h}}_{n}={\bm{y}}(n)$ is the same as the subspace spanned by the $J$ eigenvectors of the sample covariance matrix $\hat{{\bm{C}}}^{(p)}_{{\bm{y}}\mid{\bm{H}}}=\frac{1}{J}{\bm{Y}}_{p}{\bm{Y}}_{p}^{\mathrm{H}}$, which ignores the additional phase information contained in the pilot observation. Thus, the overall subspace estimate $\hat{{\bm{V}}}=[{\bm{v}}_{1},\dots,{\bm{v}}_{J}]$ is found by taking the $J$ dominant eigenvectors ${\bm{v}}_{j}$ of the sample covariance matrix defined as
$\hat{{\bm{C}}}_{{\bm{y}}\mid{\bm{H}}}=\frac{1}{N}{\bm{Y}}{\bm{Y}}^{\mathrm{H}}.$ (35)
This result has also been used in [26].
To save computational complexity while calculating the sample covariance matrix in (35), information from the previous coherence intervals can be utilized. This can be done by adaptively updating the subspace using efficient tracking algorithms as proposed in, e.g., [41, 42].
(a) Perfect Subspace Knowledge
(b) Subspace Estimation based on $N=200$ Snapshots
@@ III-D Maximum Likelihood Subspace Estimation
Figure 1: MSE over the SNR for given channel estimations based on perfect statistical knowledge in a $J=8$ user and $M=64$ antennas scenario with uncorrelated Rayleigh fading and Gaussian symbols using (a) perfect subspace knowledge and (b) $N=200$ snapshots for subspace estimation.
III-E Performance Analysis for Estimated Subspace
Fig. 1b shows the individual channel estimators’ performances based on an estimated subspace using $N=200$ snapshots for uncorrelated Rayleigh fading. Additionally, the unbiased CRBs derived in [15] and their extensions to the Bayesian Cramer-Rao bounds (BCRBs), derived in Section -C, are shown. For increasing SNR, all semi-blind channel estimators considered in this section converge to the CRBs and BCRBs. Furthermore, we see that for low SNR values the introduced MMSE-based semi-blind channel estimators outperform the different unbiased CRBs, which are lower bounds for, e.g., the EM-based estimator from [15].
In order to quantify the error introduced by the subspace estimation in comparison to perfect subspace knowledge, let us consider the projected LMMSE as introduced in Section III-B in combination with the subspace estimation from Section III-D. We assume Gaussian distributed symbols for the following derivations. The introduced error by estimating the subspace can be bounded using Theorem 2.
Theorem 2.
Let the subspace spanned by the left singular vectors of the channel matrix $\bm{H}$} be estimated based on $N$ observations. The probability that the difference between the projected LMMSE channel estimators utilizing the estimated and true subspace, respectively, is less than $\varepsilon$ is given as
$P\left(\mathbb{E}_{{\bm{X}},{\bm{N}}}[\|\hat{{\bm{h}}}_{\mathrm{proj}}(\hat{{\bm{V}}})-\hat{{\bm{h}}}_{\mathrm{proj}}({{\bm{V}}})\|^{2}]\leq\varepsilon\right)\\ \geq 1-\frac{4k^{2}(M-J)\lambda_{\max}({\bm{W}}_{\mathrm{proj}}^{\mathrm{H}}{\bm{W}}_{\mathrm{proj}})}{J^{2}N\varepsilon}\sum_{j\leq J}\frac{|s_{\max}|^{2}+2\sigma^{2}}{\lambda_{j}^{2}},$ (36)
where $\lambda_{j}$ is the $j$-th eigenvalue of ${\bm{H}}{\bm{H}}^{\mathrm{H}}$, $s_{\max}=\max_{j\in\{1,\dots,J\}}s_{j}$ with ${\bm{h}}={\bm{V}}{\bm{s}}$, and
$k^{2}=\sigma^{2}\left(\sigma^{2}+\mathrm{tr}({\bm{C}}_{{\bm{y}}\mid{\bm{H}}})\right).$ (37)
Proof.
See Appendix -D
The probability bound in Theorem 2 is very loose, as the used inequalities are not tight. Nevertheless, this bound can be used to show the superiority of the projected LMMSE based on an estimated subspace for a high number of observations compared to the other considered estimators. Table I gives the bound values for the case of $N=10^{4}$ observations at SNR =0=0 dB for different number of users. As one can see, the projected LMMSE outperforms the other estimators with high probability.
TABLE I: Lower bound on the probability of the projected LMMSE outperforming the other estimators for $N=10^{4}$ observations at SNR =0=0dB.
Number of users $J$ 88 $16$ $32$
$P(\mathrm{MSE}^{\mathrm{proj}}_{\mathrm{iid}}\leq\mathrm{MSE}^{\mathrm{sub}}_{\mathrm{iid}})$ $\geq 0.9873$ $\geq 0.9791$ $\geq 0.9642$
$P(\mathrm{MSE}^{\mathrm{proj}}_{\mathrm{iid}}\leq\mathrm{MSE}^{\mathrm{ML}}_{\mathrm{iid}})$ $\geq 0.8453$ $\geq 0.9539$ $\geq 0.9907$
$P(\mathrm{MSE}^{\mathrm{proj}}_{\mathrm{iid}}\leq\mathrm{MSE}^{\mathrm{plain}}_{\mathrm{iid}})$ $\geq 0.9946$ $\geq 0.9923$ $\geq 0.9907$
@@ IV Proposed Semi-Blind Channel Estimation - Utilizing Generative Prior
In practice, the user channels do not follow a Gaussian distribution. Hence, (10) can not be utilized directly and the CME given the pilot observation ${\bm{y}}_{p}$ is formulated as
@@
As seen in (38), the CME generally can not be computed analytically. First, the CME needs access to $p({\bm{h}})$, which is generally unavailable in practice. Additionally, no closed-form solution exists to the integral in (38).
@@
denotes the LMMSE estimate given $\bm{c}$}. However, finding a suitable condition $\bm{c}$} can be challenging, particularly as the true distribution of $\bm{h}$} is unknown. To this end, CGLMs were proposed in [30, 31, 32], which approximate the CME based on a GMM, mixtures of factor analyzers (MFA), and VAE, respectively. All three methods learn a model that provides the conditional Gaussian distribution ${\bm{h}}\mid{\bm{c}}$ based on a discrete (GMM, MFA) or continuous (VAE) latent variable $\bm{c}$}. This work focuses on the GMM and VAE, which we adapt to semi-blind channel estimation in the following.